what is 10y-13=3y+8
Found 2 solutions by elima, Bluidangel:
Answer by elima(1433)

10y-13=3y+8
We first will +13 to both sides;
10y-13+13=3y+8+13
Now we will subtract 3y from both sides;
10y-3y=3y+21-3y
Now we have the variables on one side:
7y=21
Divide both sides by 7
7y/7=21/7
y=3
Hope you understand =)

Answer by Bluidangel(2)

10y-13=3y+8
well first take a number that you think might work.
my first number was 2 so I checked my answer like this
10(2)-13=3(2)+8
20-13=6+8
7=14
so thats wrong then I took three
10(3)-13=3(3)+8
30-13=9+8
17=17 so that is correct so the answer would be
y=3